1.) Black hair in rabbits is determined by a dominant allele, B, and white hair by homozygosity for a recessive allele, b. Two heterozygous rabbits mate and produce a litter of five offspring. With what frequency would we expect to find:

A.) two whites and three black offspring

B.) four white and one black offspring.

2.) In families of size 7 offspring, what is the probability of finding the third offspring to be male and the oldest to be female?

3.) A heterozygous genetic condition called "creeper" in chickens produces shortened and deformed legs and wings, giving the bird a squatty appearance. Matings between creepers produce 775 creepers : 388 normal progeny.

A.) Is the genetic hypothesis of a 3 creeper: 1 norma ratio acceptable? Use a chi-squared test to support and explain your answer. Make sure to state the following: genetic hypothesis, null hypothesis, work and calculations to calculate X2C, the critical value: X^2T,P,df, and conclusions of the results.  

B.) Does a 2 creeper: 1 normal ratio fir the data better? Use a chi-squared test to support and explain your answer.
Answer #1

Answer

1.

In mono-hybrid cross (F2 generation show 3/4 outcome for dominant allele and 1/4 outcome for recessive allele out of 4 as when 2 heterozygous mate)

here rabbits produces 5 offspring

Bb x Bb so, according the monohybrid cross the Probability of Black hair = 3/4 and Probability of white hair = 1/4

lets solve this problem by using binomial theorem,where x is the number of black hair offspring out of 5

f(x black hair offspring) = 5Cx (3/4)x(1/4)5-x

(a) Now we have to find the probability of

Pr(Two whites and three black offspring) = 5C3 (3/4)3(1/4)2 = 0.2637

(b Pr(Four white and one black offspring) = 5C1 (3/4)1(1/4)4 = 0.0146

2.

In family, 1st offspring is female, 2nd one can be female or male, 3rd one is male, the next four off-springs will either

female or male.

Then, possible number of combinations will be = 1 * 2 * 1 * 2 * 2 * 2 * 2= 32

whereas, all possible combinations will be =2 *2 * 2 * 2 * 2 * 2 * 2= 128.

Hence, probability of finding the third offspring to be male is= 32/128 = 1/4 = 0.25.